# Learning the Boundary of Solvability: Aligning LLMs to Detect Unsolvable Problems

### v2

@@ 5 Experiments
Figure 4: Detailed Ablation Metrics. Breakdown of Solvable Accuracy (S, Red), Unsolvable Detection Rate (U, Blue), and Mean Performance (M, Green) across different configurations. The w/o UnsData w/ P setting shows a catastrophic drop in U for the 1.7B model (4.2%), indicating capability collapse. Notably, w/ UnsData w/o P achieves the highest Mean (M) score by maximizing U. The full method (w/ UnsData w/ P) demonstrates a valuable trade-off, boosting S while maintaining a high U, optimizing for practical scenarios.
Figure 5: Mechanism Analysis: Trade-off between $S$ and $U$. Visualizing the interaction between Data and Penalty. The Purple Dashed Arrow highlights the comprehensive enhancement: compared to the naive baseline (Red Square), our full method (Blue Circle) achieves significant improvements in both Solvable Accuracy ($S$) and Unsolvable Detection ($U$). Specific dynamics include: (left) Capability Collapse: Applying the penalty without data causes the 1.7B model to collapse (Red Arrow). (right) Strategic Trade-off: With adequate data, the penalty acts as a regularizer (Blue Arrow), trading a small amount of $U$ for a gain in $S$ to optimize for real-world applications.
We trained and evaluated Qwen3-1.7B and Qwen3-4B (yang2025qwen3technicalreport) in think mode on UnsolvableQA. For evaluation, we report average@8 on the AIME partitions and average@4 on other datasets. We standardize training steps: ablation variants are trained for 120 steps, while the final models are trained for 320 steps.
5.1 Main Results: Bridging the Gap with SOTA
Table 2 compares our Qwen3-UnsolRL models against both proprietary SOTA models and base baselines. The complete experimental results are presented in Appendix 4. High-reasoning LLMs (e.g., Deepseek-V3.2-R1) inherently perform well, whereas base models struggle significantly (e.g., Qwen3-1.7B Instruct achieves only 33.2% mean score). Our method bridges this gap: Qwen3-4B + UnsolRL achieves a global mean score of 80.1%, significantly outperforming GPT-5.1-low (56.7%) and approaching the performance of Gemini-3 (91.7%). Notably, our 4B model achieves the highest global Unsolvable Detection rate (90.7%), surpassing even the much larger Deepseek and Gemini models. This confirms that equipping models with the ability to identify unsolvability ($\mathcal{U}$) strengthens the rigor required to solve feasible problems ($\mathcal{S}$).
@@ 5.2 Ablation Study: Disentangling Data and Penalty
As shown in Figure 4, the naive baseline trained without unsolvable data (w/o UnsData) lacks the fundamental capability to recognize unsolvability. For Qwen3-4B, detection is weak ($U=64.1\%$). Introducing explicit unsolvable data provides the necessary supervision, boosting detection significantly ($U\to 89.3\%$). This confirms that instruction following alone is insufficient; explicit data is the foundational requirement for boundary learning.
@@
Figure 5(left) illustrates a critical failure mode caused by objective mismatch. When the penalty is enforced without providing ground-truth unsolvable examples, the model, lacking positive examples of valid refusal, minimizes the penalty risk by defaulting to answering all queries. This behavior effectively collapses its refusal mechanism, as seen in the 1.7B model’s detection rate plummeting from 70.8% to 4.2%.
@@
When grounded with adequate data, the penalty shifts from being a suppressor to a regularizer. For Qwen3-4B, adding the penalty (w/ UnsData w/ P) improves Solvable Accuracy ($S$) from 60.6% to 62.6%, at the cost of a slight drop in detection ($U:89.3\%\to 84.8\%$). This trade-off is strategic: it sacrifices marginal detection performance to enforce stricter reasoning rigor on feasible problems, optimizing the model for real-world utility.
@@
Viewing the trajectory globally (the purple dashed arrow in Figure 5b), our full method represents a comprehensive upgrade over the naive baseline (w/o UnsData w/o P). We achieve significant gains in both dimensions for Qwen3-4B ($S:57.7\%\to 62.6\%$, $U:64.1\%\to 84.8\%$). This confirms that UnsolvableRL does not merely shift the decision boundary, but fundamentally enhances both reliability and reasoning capability.
@@ 5.3 Comparative Analysis: TruthRL and Calibration
We further benchmark our method against TruthRL(wei2025truthrl) and different thresholding strategies. The results are summarized in Figure 6.
@@
As shown in the first two panels of Figure 6, TruthRL tends to be overly conservative, leading to over-refusal. For Qwen3-1.7B, while TruthRL achieves a high Rejection Rate (18.58%), its Accuracy drops to 40.4%, which is inferior to even the baseline. In contrast, our method achieves a significantly higher Accuracy (50.3%) with a moderate and precise Rejection Rate (7.96%). This indicates that our False Unsolvability Penalty provides a softer, more distinctive signal than TruthRL’s hard refusal reward, preserving the model’s core reasoning capabilities.
@@
The rightmost panel of Figure 6 highlights the importance of our dynamic calibration. The Fixed-$\tau$ baseline yields a low rejection rate (4.36%), suggesting overconfidence. By dynamically adjusting $\tau$ based on the moving average of solvable confidence, our method (w/ P Final) maintains a robust rejection rate (6.87%) without sacrificing accuracy (80.1%), ensuring the model remains prudent even as it becomes stronger.
@@ 6 Conclusion
In this paper, we address the challenge of LLM reliability by formally decoupling objective unsolvability from subjective capability. We introduce UnsolvableQA for rigorous evaluation and UnsolvableRL, an alignment framework featuring dynamic calibration. Empirically, our method achieves a Pareto improvement over baselines: it boosts solvable accuracy (Overall $S$) from 43.4% to 69.4% while maintaining robust unsolvability detection ($U>90\%$ on Qwen3-4B). Crucially, our analysis uncovers the dual nature of alignment penalties: without grounding data, they trigger Capability Collapse; yet with adequate data, they facilitate a Strategic Trade-off, acting as a Regularizer for Rigor that prevents hallucination without the over-refusal pitfalls observed in prior methods like TruthRL. We hope this work establishes a new paradigm for building models that rigorously distinguish between inherent unsolvability and problems that simply exceed their current capabilities.
@@
While the UnsolvableQA dataset is relatively comprehensive within the selected domains (logic puzzles, mathematics, mazes), it may not fully encompass the various forms of inherently unsolvable problems encountered in real-world applications. Future work could expand the dataset to include more diverse and ambiguous types of unsolvable problems.
@@ A.1 Comprehensive experimental setup and results
Deepseek-V3.2-R 98.0 77.5 87.8 83.5 94.0 88.8 84.0 96.0 90.0 70.0 86.0 78.0 100.0 98.9 99.5 98.0 98.0 98.0 85.0 63.8 74.4 88.4 87.7 88.1
GPT-5.1-Low 14.0 50.0 32.0 8.3 82.0 45.1 50.0 86.0 68.0 6.0 12.0 9.0 95.0 100.0 97.5 86.0 94.0 90.0 61.7 48.9 55.3 45.9 67.6 56.7
Gemini-3.0-Pro 94.0 94.0 94.0 100.0 98.0 99.0 100.0 90.0 95.0 98.0 100.0 99.0 99.0 94.6 96.8 96.0 91.0 93.5 95.0 34.1 64.5 97.4 86.0 91.7
Qwen3-4B-Instruct 81.0 17.0 49.0 37.5 57.0 47.3 37.0 56.5 46.8 34.5 6.5 20.5 32.7 55.6 44.1 13.3 64.5 38.9 67.9 25.3 46.6 43.4 40.3 41.9
+ TruthRL 91.0 91.0 91.0 32.3 86.0 59.2 49.0 90.0 69.5 38.0 15.0 26.5 90.0 77.1 83.6 31.0 56.5 43.8 67.5 25.5 46.5 57.0 63.0 60.0
+ w/o UnsData w/o P 83.0 92.0 87.5 36.5 81.0 58.8 47.0 84.0 65.5 46.0 20.0 33.0 88.5 76.6 82.6 34.5 60.5 47.5 68.3 34.6 51.5 57.7 64.1 60.9
+ w/o UnsData w/ P 90.5 12.5 51.5 39.6 67.5 53.5 49.5 64.5 57.0 56.5 4.0 30.3 94.8 83.2 89.0 56.3 72.0 64.1 69.0 18.4 43.7 65.2 46.0 55.6
+ w/ UnsData w/o P 84.0 95.0 89.5 39.6 96.0 67.8 53.0 100.0 76.5 45.0 87.0 66.0 95.5 95.2 95.4 40.0 89.5 64.8 67.1 62.2 64.7 60.6 89.3 75.0
+ w/ UnsData w/ P 88.0 97.0 92.5 35.9 95.5 65.7 51.5 99.5 75.5 48.5 71.0 59.8 94.0 95.5 94.8 53.3 79.8 66.6 66.9 55.3 61.1 62.6 84.8 73.7
+ Fixed-$\tau$ Final 87.5 99.0 93.3 39.1 100.0 69.6 67.0 99.5 83.3 51.0 98.0 74.5 96.3 100.0 98.2 52.5 95.8 74.2 69.0 63.6 66.3 66.1 93.7 79.9
+ w/ UnsData w/ P Final 94.5 99.0 96.8 41.1 94.5 67.8 55.0 96.5 75.8 63.5 94.5 79.0 96.5 98.9 97.7 65.8 89.0 77.9 69.6 62.2 65.9 69.4 90.7 80.1
Qwen3-1.7B-Instruct 78.0 23.0 50.5 22.9 28.0 25.5 14.0 33.0 23.5 8.0 21.0 14.5 0.0 85.6 42.8 0.0 80.0 40.0 38.3 33.5 35.9 23.0 43.4 33.2
+ TruthRL 76.0 80.0 78.0 19.8 49.0 34.4 17.0 58.0 37.5 4.0 25.0 14.5 1.5 82.4 42.0 0.0 73.5 36.8 42.1 37.2 39.7 22.9 57.9 40.4
+ w/o UnsData w/o P 77.0 88.0 82.5 25.0 60.0 42.5 15.0 83.0 49.0 5.0 53.0 29.0 0.5 93.6 47.1 0.0 79.5 39.8 32.9 38.3 35.6 22.2 70.8 46.5
@@
+ w/ UnsData w/o P 70.0 97.0 83.5 22.9 94.0 58.5 22.0 97.0 59.5 4.0 79.0 41.5 0.5 95.2 47.9 0.0 86.0 43.0 31.7 56.4 44.0 21.6 86.4 54.0
+ w/ UnsData w/ P 82.0 96.5 89.3 18.2 85.0 51.6 20.5 89.5 55.0 8.0 62.0 35.0 0.0 90.7 45.4 0.0 74.3 37.2 40.8 36.4 38.6 24.2 76.3 50.3
+ w/ UnsData w/ P Final 92.0 0.50 92.5 54.4 32.19 86.6 56.0 21.50 77.5 35.0 2.00 37.0 47.6 0.55 48.2 39.5 6.75 46.3 44.1 3.17 47.3 52.7 9.52 62.2
@@
Deepseek-V3.2-R 87.8 0.00 87.8 88.8 8.30 97.1 90.0 6.00 96.0 78.0 9.00 87.0 99.5 0.00 99.5 98.0 0.00 98.0 74.4 0.80 75.2 88.1 3.44 91.5
GPT-5.1-Low 32.0 1.00 33.0 45.1 2.08 47.2 68.0 8.00 76.0 9.0 1.00 10.0 97.5 0.00 97.5 90.0 0.00 90.0 55.3 7.90 63.2 56.7 2.85 59.6
Gemini-3.0-Pro 94.0 1.00 95.0 99.0 0.00 99.0 95.0 0.00 95.0 99.0 0.00 99.0 96.8 0.00 96.8 93.5 0.00 93.5 64.5 0.00 64.5 91.7 0.14 91.8
Qwen3-4B-Instruct 49.0 0.00 49.0 47.3 24.74 72.0 46.8 12.00 58.8 20.5 4.00 24.5 44.1 0.00 44.1 38.9 0.25 39.2 46.6 0.00 46.6 41.9 5.86 47.7
+ TruthRL 91.0 0.50 91.5 59.2 34.15 93.4 69.5 16.00 85.5 26.5 18.50 45.0 83.6 2.50 86.1 43.8 11.00 54.8 46.5 0.00 46.5 60.0 11.81 71.8
+ w/o UnsData w/o P 87.5 0.00 87.5 58.8 28.85 87.7 65.5 15.50 81.0 33.0 11.00 44.0 82.6 0.00 82.6 47.5 0.00 47.5 51.5 0.00 51.5 60.9 7.91 68.8
+ w/o UnsData w/ P 51.5 0.00 51.5 53.5 21.09 74.6 57.0 7.00 64.0 30.3 2.25 32.6 89.0 0.00 89.0 64.1 0.13 64.2 43.7 0.00 43.7 55.6 4.35 60.0
+ w/ UnsData w/o P 89.5 1.50 91.0 67.8 19.00 86.8 76.5 10.00 86.5 66.0 13.00 79.0 95.4 1.00 96.4 64.8 4.00 68.8 64.7 0.42 65.1 75.0 7.00 82.0
+ w/ UnsData w/ P 92.5 1.00 93.5 65.7 19.75 85.5 75.5 15.50 91.0 59.8 7.50 67.3 94.8 0.00 94.8 66.6 1.00 67.6 61.1 0.00 61.1 73.7 6.39 80.1
+ Fixed-$\tau$ Final 93.3 0.00 93.3 69.6 18.49 88.1 83.3 12.00 95.3 74.5 0.00 74.5 98.2 0.00 98.2 74.2 0.00 74.2 66.3 0.00 66.3 79.9 4.36 84.3
+ w/ UnsData w/ P Final 96.8 0.00 96.8 67.8 26.82 94.6 75.8 19.00 94.8 79.0 2.25 81.3 97.7 0.00 97.7 77.1 0.00 77.1 65.9 0.00 65.9 80.1 6.87 87.0
Qwen3-1.7B-Instruct 50.5 0.00 50.5 25.5 8.85 34.3 23.5 12.00 35.5 14.5 0.50 15.0 42.8 0.00 42.8 40.0 0.00 40.0 35.9 0.00 35.9 33.2 3.05 36.3
+ TruthRL 78.0 2.00 80.0 34.4 29.58 64.0 37.5 39.50 77.0 14.5 16.50 31.0 42.0 12.78 54.8 36.8 24.00 60.8 39.7 5.68 45.4 40.4 18.58 59.0
+ w/o UnsData w/o P 82.5 0.50 83.0 42.5 29.17 71.7 49.0 30.00 79.0 29.0 6.50 35.5 47.1 0.25 47.3 39.8 3.50 43.3 35.6 4.17 39.8 46.5 10.58 57.1
+ w/o UnsData w/ P 43.5 0.25 43.8 16.5 19.01 35.5 19.8 8.00 27.8 7.3 5.25 12.5 0.0 0.00 0.0 0.0 0.00 0.0 20.7 0.00 20.7 15.4 4.64 20.1
+ w/ UnsData w/o P 83.5 1.00 84.5 58.5 22.88 81.4 59.5 19.50 79.0 41.5 2.50 44.0 47.9 1.50 49.4 43.0 0.75 43.8 44.0 2.52 46.5 54.0 7.24 61.2
+ w/ UnsData w/ P 89.3 0.00 89.3 51.6 27.77 79.4 55.0 19.50 74.5 35.0 2.50 37.5 45.4 0.00 45.4 37.2 4.50 41.7 38.6 1.46 40.1 50.3 7.96 58.3
+ w/ UnsData w/ P Final 92.0 0.50 92.5 54.4 32.19 86.6 56.0 21.50 77.5 35.0 2.00 37.0 47.6 0.55 48.2 39.5 6.75 46.3 44.1 3.17 47.3 52.7 9.52 62.2

### v3

@@ 5 Experiments
Figure 5: Visualizing the interaction between Solvable Accuracy ($S$) and Unsolvable Detection ($U$). The Purple Arrow highlights our full method (Blue Circle) achieving a Pareto improvement over the baseline (Red Square). Both models exhibit consistent dynamics: applying penalties without Unsolvable Data (UD) triggers Capability Collapse (Red Arrow), sacrificing $U$ to maximize $S$. Conversely, when paired with UD, the penalty acts as a Regularizer (Blue Arrow), trading marginal $U$ for significant gains in $S$ compared to using UD alone (Blue Square).
We trained and evaluated Qwen3-1.7B and Qwen3-4B (Yang et al., 2025) in think mode on UnsolvableQA. For evaluation, we report average@8 on the AIME partitions and average@4 on other datasets. We standardize training steps: ablation variants are trained for 120 steps, while the final models are trained for 320 steps.
5.1 Main Results: Comparative Analysis with SOTA Models
Table 2 compares our Qwen3-UnsolvableRL models against both proprietary SOTA models and base baselines. The complete experimental results are presented in Appendix 4, which also verifies the generalizability of our method on Llama models, where UnsolvableRL boosts the overall Mean score to 55.1% compared to the 25.2% baseline. High-reasoning LLMs (e.g., Deepseek-V3.2-R1) inherently perform well, whereas base models struggle significantly (e.g., Qwen3-1.7B Instruct achieves only 32.4% mean score). Our method bridges this gap: Qwen3-4B + UnsolvableRL achieves a global mean score of 78.6%, significantly outperforming GPT-5.1-low (56.2%) and approaching Gemini-3 (90.8%). Notably, our 4B model achieves the highest global Unsolvable Detection rate (87.5%), surpassing even the much larger Deepseek and Gemini models. This confirms that equipping models with the ability to identify unsolvability ($\mathcal{U}$) strengthens the rigor required to solve feasible problems ($\mathcal{S}$). We further validated our method on two OOD benchmarks in Appendix A.2.
@@ 5.2 Ablation Study: Disentangling Data and Penalty
As shown in Figure 4, the naive baseline trained without unsolvable data (w/o UnsData) lacks the fundamental capability to recognize unsolvability. For Qwen3-4B, detection is weak ($U=62.6\%$). Introducing explicit unsolvable data provides the necessary supervision, boosting detection significantly ($U\to 84.8\%$). This confirms that optimizing solely for reasoning correctness on solvable tasks is insufficient.
@@
Figure 5(left) illustrates a critical failure mode caused by objective mismatch. When the penalty is enforced without providing ground-truth unsolvable examples, the model, lacking positive examples of valid refusal, minimizes the penalty risk by defaulting to answering all queries. This behavior effectively collapses its refusal mechanism, as seen in the 1.7B model’s detection rate plummeting from 68.8% to 4.2%.
@@
When grounded with adequate data, the penalty shifts from being a suppressor to a regularizer. For Qwen3-4B, adding the penalty (w/ UnsData w/ P) improves Solvable Accuracy ($S$) from 60.6% to 62.6%, at the cost of a slight drop in detection ($U:84.8\%\to 81.5\%$). This trade-off is strategic: it sacrifices marginal detection performance to enforce stricter reasoning rigor on feasible problems, optimizing the model for real-world utility.
@@
Viewing the trajectory globally (the purple dashed arrow in Figure 5b), our full method represents a comprehensive upgrade over the naive baseline (w/o UnsData w/o P). We achieve significant gains in both dimensions for Qwen3-4B ($S:57.7\%\to 62.6\%$, $U:62.6\%\to 81.5\%$). This confirms that UnsolvableRL does not merely shift the decision boundary, but fundamentally enhances both reliability and reasoning capability.
@@ 5.3 Comparative Analysis: TruthRL and Calibration
We further benchmark our method against TruthRL (Wei et al., 2025) and different thresholding strategies.
@@
As shown in the first two panels of Figure 6, TruthRL tends to be overly conservative, leading to over-refusal. For Qwen3-1.7B, while TruthRL achieves a high Rejection Rate (18.58%), its Accuracy drops to 39.0%, which is inferior to even the baseline. In contrast, our method achieves a significantly higher Accuracy (45.5%) with a moderate and precise Rejection Rate (10.58%). This indicates that our target accuracy threshold provides a softer, more distinctive signal than TruthRL’s hard refusal reward, preserving the model’s core reasoning capabilities.
@@
The rightmost panel of Figure 6 highlights the importance of our dynamic calibration. The Fixed-$\tau$ baseline yields a low rejection rate (4.36%), suggesting overconfidence. By dynamically adjusting $\tau$ based on the moving average of solvable confidence, our method (w/ P Final) maintains a robust rejection rate (6.87%) without sacrificing accuracy (78.6%), ensuring the model remains prudent even as it becomes stronger.
@@ 6 Conclusion
In this paper, we address the challenge of LLM reliability by formally decoupling objective unsolvability from subjective capability. We introduce UnsolvableQA for rigorous evaluation and UnsolvableRL, an alignment framework featuring dynamic calibration. Empirically, our method achieves a Pareto improvement over baselines: it boosts solvable accuracy (Overall $S$) from 43.4% to 69.4% while maintaining robust unsolvability detection ($U>85\%$ on Qwen3-4B). Crucially, our analysis uncovers the dual nature of alignment penalties: without grounding data, they trigger Capability Collapse; yet with adequate data, they facilitate a Strategic Trade-off, acting as a Regularizer for Rigor that prevents hallucination without the over-refusal pitfalls observed in prior methods like TruthRL. We hope this work establishes a new paradigm for building models that rigorously distinguish between inherent unsolvability and problems that simply exceed their current capabilities.
@@
While the UnsolvableQA dataset is relatively comprehensive within the selected domains (logic puzzles, mathematics, mazes), it may not fully encompass the various forms of inherently unsolvable problems encountered in real-world applications. Future work could expand the dataset to include more diverse and ambiguous types of unsolvable problems. Additionally, due to computational resource constraints, we were unable to conduct a more exhaustive set of ablation studies, such as scaling experiments on larger model architectures.
@@ A.1 Comprehensive experimental setup and results
Deepseek-V3.2-R 98.0 77.5 87.8 83.5 94.0 88.8 84.0 96.0 90.0 70.0 86.0 78.0 100.0 98.9 99.5 98.0 98.0 98.0 85.0 40.3 62.7 88.4 84.4 86.5
GPT-5.1-Low 14.0 50.0 32.0 8.3 82.0 45.1 50.0 86.0 68.0 6.0 12.0 9.0 95.0 100.0 97.5 86.0 94.0 90.0 61.7 42.5 52.1 45.9 66.6 56.2
Gemini-3.0-Pro 94.0 94.0 94.0 100.0 98.0 99.0 100.0 90.0 95.0 98.0 100.0 99.0 99.0 94.6 96.8 96.0 91.0 93.5 95.0 21.2 58.1 97.4 84.1 90.8
Qwen3-4B-Instruct 81.0 17.0 49.0 37.5 57.0 47.3 37.0 56.5 46.8 34.5 6.5 20.5 32.7 55.6 44.1 13.3 64.5 38.9 67.9 14.8 41.4 43.4 38.8 41.1
+ TruthRL 91.0 91.0 91.0 32.3 86.0 59.2 49.0 90.0 69.5 38.0 15.0 26.5 90.0 77.1 83.6 31.0 56.5 43.8 67.5 19.1 43.3 57.0 62.1 59.6
+ w/o UnsData w/o P 83.0 92.0 87.5 36.5 81.0 58.8 47.0 84.0 65.5 46.0 20.0 33.0 88.5 76.6 82.6 34.5 60.5 47.5 68.3 24.0 46.2 57.7 62.6 60.2
+ w/o UnsData w/ P 90.5 12.5 51.5 39.6 67.5 53.5 49.5 64.5 57.0 56.5 4.0 30.3 94.8 83.2 89.0 56.3 72.0 64.1 69.0 7.4 38.2 65.2 44.4 54.8
+ w/ UnsData w/o P 84.0 95.0 89.5 39.6 96.0 67.8 53.0 100.0 76.5 45.0 87.0 66.0 95.5 95.2 95.4 40.0 89.5 64.8 67.1 30.8 49.0 60.6 84.8 72.7
+ w/ UnsData w/ P 88.0 97.0 92.5 35.9 95.5 65.7 51.5 99.5 75.5 48.5 71.0 59.8 94.0 95.5 94.8 53.3 79.8 66.6 66.9 31.9 49.4 62.6 81.5 72.0
+ Fixed-$\tau$ Final 87.5 99.0 93.3 39.1 100.0 69.6 67.0 99.5 83.3 51.0 98.0 74.5 96.3 100.0 98.2 52.5 95.8 74.2 69.0 41.9 55.5 66.1 90.6 78.4
+ w/ UnsData w/ P Final 94.5 99.0 96.8 41.1 94.5 67.8 55.0 96.5 75.8 63.5 94.5 79.0 96.5 98.9 97.7 65.8 89.0 77.9 69.6 40.4 55.0 69.4 87.5 78.6
Qwen3-1.7B-Instruct 78.0 23.0 50.5 22.9 28.0 25.5 14.0 33.0 23.5 8.0 21.0 14.5 0.0 85.6 42.8 0.0 80.0 40.0 38.3 21.2 29.8 23.0 41.7 32.4
+ TruthRL 76.0 80.0 78.0 19.8 49.0 34.4 17.0 58.0 37.5 4.0 25.0 14.5 1.5 82.4 42.0 0.0 73.5 36.8 42.1 17.0 29.6 22.9 55.0 39.0
+ w/o UnsData w/o P 77.0 88.0 82.5 25.0 60.0 42.5 15.0 83.0 49.0 5.0 53.0 29.0 0.5 93.6 47.1 0.0 79.5 39.8 32.9 24.4 28.7 22.2 68.8 45.5
@@
+ w/ UnsData w/o P 70.0 97.0 83.5 22.9 94.0 58.5 22.0 97.0 59.5 4.0 79.0 41.5 0.5 95.2 47.9 0.0 86.0 43.0 31.7 46.8 39.3 21.6 85.0 53.3
+ w/ UnsData w/ P 82.0 96.5 89.3 18.2 85.0 51.6 20.5 89.5 55.0 8.0 62.0 35.0 0.0 90.7 45.4 0.0 74.3 37.2 40.8 21.2 31.0 24.2 74.2 49.2
+ w/ UnsData w/ P Final 84.0 100.0 92.0 20.8 88.0 54.4 27.0 85.0 56.0 11.0 59.0 35.0 0.0 95.2 47.6 0.0 79.0 39.5 35.4 28.7 32.1 25.5 76.4 50.9
Llama3-8B-Distill 28.0 48.0 38.0 27.1 2.0 14.6 14.0 0.0 7.0 6.0 56.0 31.0 1.0 44.7 22.9 0.0 52.0 26.0 35.0 39.4 37.2 15.9 34.6 25.2
+ w/o UnsData w/o P 58.0 2.0 30.0 41.7 0.0 20.8 50.0 0.0 25.0 6.0 2.0 4.0 67.0 7.4 37.2 8.0 8.0 8.0 46.7 9.6 28.1 39.6 4.1 21.9
+ w/ UnsData w/ P 66.0 4.0 35.0 39.6 98.0 68.8 48.0 90.0 69.0 10.0 78.0 44.0 60.0 3.2 31.6 7.0 20.0 13.5 50.0 30.9 40.5 40.1 46.3 43.2
+ w/ UnsData w/ P Final 80.0 22.0 51.0 52.1 100.0 76.1 74.0 92.0 83.0 30.0 76.0 53.0 67.0 42.6 54.8 13.0 35.0 24.0 43.3 44.7 44.0 51.3 58.9 55.1
@@
Deepseek-V3.2-R 87.8 0.00 87.8 88.8 8.30 97.1 90.0 6.00 96.0 78.0 9.00 87.0 99.5 0.00 99.5 98.0 0.00 98.0 62.7 0.80 63.5 86.5 3.44 90.0
GPT-5.1-Low 32.0 1.00 33.0 45.1 2.08 47.2 68.0 8.00 76.0 9.0 1.00 10.0 97.5 0.00 97.5 90.0 0.00 90.0 52.1 7.90 60.0 56.2 2.85 59.1
Gemini-3.0-Pro 94.0 1.00 95.0 99.0 0.00 99.0 95.0 0.00 95.0 99.0 0.00 99.0 96.8 0.00 96.8 93.5 0.00 93.5 58.1 0.00 58.1 90.8 0.14 90.9
Qwen3-4B-Instruct 49.0 0.00 49.0 47.3 24.74 72.0 46.8 12.00 58.8 20.5 4.00 24.5 44.1 0.00 44.1 38.9 0.25 39.2 41.4 0.00 41.4 41.1 5.86 47.0
+ TruthRL 91.0 0.50 91.5 59.2 34.15 93.4 69.5 16.00 85.5 26.5 18.50 45.0 83.6 2.50 86.1 43.8 11.00 54.8 43.3 0.00 43.3 59.6 11.81 71.4
+ w/o UnsData w/o P 87.5 0.00 87.5 58.8 28.85 87.7 65.5 15.50 81.0 33.0 11.00 44.0 82.6 0.00 82.6 47.5 0.00 47.5 46.2 0.00 46.2 60.2 7.91 68.1
+ w/o UnsData w/ P 51.5 0.00 51.5 53.5 21.09 74.6 57.0 7.00 64.0 30.3 2.25 32.6 89.0 0.00 89.0 64.1 0.13 64.2 38.2 0.00 38.2 54.8 4.35 59.2
+ w/ UnsData w/o P 89.5 1.50 91.0 67.8 19.00 86.8 76.5 10.00 86.5 66.0 13.00 79.0 95.4 1.00 96.4 64.8 4.00 68.8 49.0 0.42 49.4 72.7 7.00 79.7
+ w/ UnsData w/ P 92.5 1.00 93.5 65.7 19.75 85.5 75.5 15.50 91.0 59.8 7.50 67.3 94.8 0.00 94.8 66.6 1.00 67.6 49.4 0.00 49.4 72.0 6.39 78.4
+ Fixed-$\tau$ Final 93.3 0.00 93.3 69.6 18.49 88.1 83.3 12.00 95.3 74.5 0.00 74.5 98.2 0.00 98.2 74.2 0.00 74.2 55.5 0.00 55.5 78.4 4.36 82.8
+ w/ UnsData w/ P Final 96.8 0.00 96.8 67.8 26.82 94.6 75.8 19.00 94.8 79.0 2.25 81.3 97.7 0.00 97.7 77.1 0.00 77.1 55.0 0.00 55.0 78.6 6.87 85.5
Qwen3-1.7B-Instruct 50.5 0.00 50.5 25.5 8.85 34.3 23.5 12.00 35.5 14.5 0.50 15.0 42.8 0.00 42.8 40.0 0.00 40.0 29.8 0.00 29.8 32.4 3.05 35.5
+ TruthRL 78.0 2.00 80.0 34.4 29.58 64.0 37.5 39.50 77.0 14.5 16.50 31.0 42.0 12.78 54.8 36.8 24.00 60.8 29.6 5.68 35.3 39.0 18.58 57.6
+ w/o UnsData w/o P 82.5 0.50 83.0 42.5 29.17 71.7 49.0 30.00 79.0 29.0 6.50 35.5 47.1 0.25 47.3 39.8 3.50 43.3 28.7 4.17 32.9 45.5 10.58 56.1
+ w/o UnsData w/ P 43.5 0.25 43.8 16.5 19.01 35.5 19.8 8.00 27.8 7.3 5.25 12.5 0.0 0.00 0.0 0.0 0.00 0.0 20.7 0.00 20.7 15.4 4.64 20.0
+ w/ UnsData w/o P 83.5 1.00 84.5 58.5 22.88 81.4 59.5 19.50 79.0 41.5 2.50 44.0 47.9 1.50 49.4 43.0 0.75 43.8 39.3 2.52 41.8 53.3 7.24 60.5
+ w/ UnsData w/ P 89.3 0.00 89.3 51.6 27.77 79.4 55.0 19.50 74.5 35.0 2.50 37.5 45.4 0.00 45.4 37.2 4.50 41.7 31.0 1.46 32.5 49.2 7.96 57.2
+ w/ UnsData w/ P Final 92.0 0.50 92.5 54.4 32.19 86.6 56.0 21.50 77.5 35.0 2.00 37.0 47.6 0.55 48.2 39.5 6.75 46.3 32.1 3.17 35.3 50.9 9.52 60.4
Llama3-8B-Distill 38.0 22.00 60.0 14.6 43.75 58.4 7.0 34.00 41.0 31.0 56.00 87.0 22.9 34.19 57.1 26.0 44.50 70.5 37.2 6.78 44.0 25.2 34.46 59.7
+ w/o UnsData w/o P 30.0 0.00 30.0 20.8 5.21 26.0 25.0 4.00 29.0 4.0 6.00 10.0 37.2 1.00 38.2 8.0 4.00 12.0 28.1 0.83 28.9 21.9 3.01 24.9
+ w/ UnsData w/ P 35.0 0.00 35.0 68.8 16.67 85.5 69.0 3.00 72.0 44.0 19.00 63.0 31.6 3.00 34.6 13.5 8.50 22.0 40.5 0.83 41.3 43.2 7.29 50.5
+ w/ UnsData w/ P Final 51.0 2.00 53.0 76.1 10.42 86.5 83.0 4.00 87.0 53.0 19.00 72.0 54.8 4.05 58.9 24.0 17.00 41.0 44.0 3.34 47.3 55.1 8.54 63.7
@@
We also extended our evaluation to the Llama3-8B architecture to verify the generalizability of our method. As shown in the tables, training exclusively on solvable data (w/o UnsData w/o P) yields a dramatic improvement in solvable accuracy compared to the baseline ($S:15.9\%\to 39.6\%$). However, this comes at a catastrophic cost: the model’s ability to detect unsolvable problems collapses ($U:34.6\%\to 4.1\%$), leading to a net decline in the overall Mean score ($M:25.2\%\to 21.9\%$).
In contrast, our proposed method (w/ UnsData w/ P) successfully resolves this conflict. At step 120, it maintains superior solvable accuracy ($S=40.1\%$) while simultaneously robustifying unsolvable detection ($U=46.3\%$), nearly doubling the overall Mean score to 43.2%. Furthermore, extending the training to approximately 400 steps (Final) demonstrates robust scalability: the model achieves further gains with Solvable Accuracy reaching 51.3% and Unsolvable Detection rising to 58.9%, resulting in a comprehensive Mean score of 55.1%. This confirms that UnsolvableRL effectively prevents capability collapse and enhances reasoning rigor across different model families.